Q1. John derives utility from consuming good X and Y.
His utility function is given by U(X,Y) = LnX + 10Y.
The price of good Y is 1 and the price of good X is Px and his income is M.
(a) Derive John’s demand for X and Y.
(b) Suppose the price of X fell from $2 to $1, calculate the substitution effect and
income effect for the change in X.

Need help in part b
1
Expert's answer
2020-11-27T13:09:19-0500
"Solution"

"U(X,Y)=LnX+10Y\\
Derive\ the\ first\ order\ conditions\ for\ X\ and\ Y and\ then\ solve\ for\ Y.\\
MU_X=\frac{\Delta\ U }{\Delta\ X}=Ln\\
MU_Y=\frac{\Delta\ U}{\Delta\ Y}=10\\
Marginal\ rate\ of\ substitution(MRS)\\
=\frac{MU_Y}{MU_Y}=\frac{Ln}{10}\\
Ln=10P_x\\
M=P_XQ_X+Q_Y\\
=\frac{ln}{10}QX+Q_Y\\
Q_Y=\frac{Ln}{10}Q_X"